HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

MODELLING OF PLANETARY ATMOSPHERES FROM ACCURATE MOLECULAR LINE LISTS: PERSPECTIVES IN QUANTUM-CHEMICAL

AND NUCLEAR-MOTION CALCULATIONS

OLEG EGOROV, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State University 36, Lenin Ave., Tomsk, 634050 Russia;

\underline{MICHAËL REY} , \underline{DOMINIKA VIGLASKA} , Groupe de Spectrométrie Moléculaire et Atmosphérique UMR CNRS 7331, UFR Sciences BP 1039, 51687

Reims Cedex 2, France; ANDREI V. NIKITIN, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia, Faculty of Physics Tomsk State University 36,

Lenin Ave., Tomsk, 634050 Russia

The study of various planetary objects, including the modelling of Earth atmosphere, remains an active field of research and demonstrates how having access to accurate molecular line lists is of primary importance. In this poster, we propose a general recipe for the construction of future line lists for different families of semirigid and nonrigid molecules (including short-lived radicals) up to a dozen atoms as well as for small van der Waals (vdW) complexes for which the knowledge of intra- and intermolecular states is of great interest. Due to their constant progress, we deeply think that ab initio calculation will be inseparable from high-resolution spectroscopy in a near future, by providing complementary information to the ordinary effective perturbative models. To this end, we intend using existing quantum chemistry methods as efficiently as possible.

− In this poster, we will briefly explain how to gradually increase accuracy of the ab initio surfaces from suitable corrections and basis-set extrapolations for applications requiring (or not) sub-percent accuracy for the line intensities. With the increase in the computer capabilities, it is now reasonable to obtain ∼ 0.1 cm^{−1} of accuracy for the energy levels, and even less for the pure ab initio transitions, before applying a refinement procedure.

− In parallel, we will show how the methodology we adopted these past few years, including efficient algorithms for solving the time-independent Schrödinger equation and mathematical tools designed to spectroscopists, will be of great help for constructing suitable tailor-made Hamiltonian (polyad) models adapted to the low, medium and high resolution.

Among the molecules of planetary interest that could be potentially studied in short-, medium- and long-term, we can mention the medium-sized rigid and nonrigid systems H_{2}O_{2}, HONO, HOCO, HCOOH, HNO_{3}, CH_{3}CN, C_{2}H_{2}O, CH_{3}OH or CH_{3}NH_{2} and larger systems as C_{5}H_{6}, C_{6}H_{6} or C_{4}H_{4}O_{2}N_{2}.

A potential list of vdW molecules could include the water clusters (dimers: H_{2}O-HX, H_{2}O-O_{3}, H_{2}O-CO_{2}, H_{2}O-H_{2}O, trimers), (HF)_{3} but also other species like CH_{4}-

p-number: p103  Submitted on Tue, 13 Jun 2023 15:40:20 +0200

N_{2} or CH_{4}-O_{2}. In a general manner, nuclear motion calculations of vdW complexes beyond the rigid-monomer approximations will be considered.

This work is supported by the french Agence Nationale de la Recherche (ANR TEMMEX, Grant No. 21-CE30-0053-01) and the Russian Scientific Foundation (RSF, No. 22-42-09022).

p-number: p103  Submitted on Tue, 13 Jun 2023 15:40:20 +0200